# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v1

@@ A.1 Pruning
As in the supervised learning setting, decision trees for causal effect estimation (including in the second stage of CDT) are often pruned to avoid overfitting. There are generally two types of pruning: pre-pruning and post-pruning. Pre-pruning refers to techniques or early stopping criteria that are used to stop growing the decision tree during the tree construction process. Common pre-pruning approaches include setting constraints on the maximum depth of the tree, the minimum number of samples per node, or the minimum information gain from making each split. Post-pruning refers to techniques used to remove uninformative nodes from the decision tree after it has already been constructed. For example, in the rpart R package (Therneau et al., 1997), cross-validation is performed to choose the best complexity parameter $\alpha$, where $\alpha$ measures the cost of adding another split to the tree, and all nodes with a cost greater than $\alpha$ are subsequently removed or pruned from the tree.
Through simulations (see Section 5), we demonstrate that standard pre-pruning and post-pruning approaches used in ordinary decision trees (e.g., CART (Breiman et al., 1984)) also work well for the student decision tree model in CDT and are hence recommended. In particular, standard here means using the default pre-pruning criteria in the rpart R package and post-pruning based upon the cross-validated complexity parameter $\alpha$, as described previously. We provide additional simulation results for pruning in Appendix D.
@@ A.2 Extended discussion on valid teacher models
Let $s$ and $s^{d}$ respectively be the optimal splits using the original individual-level treatment effects $\tau_{i}$ and distilled individual-level treatment effects $\hat{\tau}_{i}^{d}$, as defined in Assumption (3.2). Also, define $v_{i}:=\tau_{i}-\hat{\tau}_{i}^{d}$ for each $i=1,\ldots,n$. We can then write
@@
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]\leq 2\left\lvert\hat{s}_{n}-s^{d}\right\rvert+\varepsilon,$
where $\varepsilon=2|s^{d}-s|$.
A.3 Non-parametric test of treatment effect heterogeneity
We propose a nonparametric test of treatment effect heterogeneity. Following Imai and Li (2024), we consider the following null hypothesis of no treatment effect heterogeneity:
$H_{0}:\tau^{(1)}=\ldots=\tau^{(G)}=\tau,$ (9)
for all subgroups $g\in\{1,\ldots,G\}$, with the following estimator:
A.4 Improving Finite-Sample Performance
While the main manuscript focused on experimental settings, we can similarly estimate the subgroup average treatment effect in observational settings. However, we must account for potential confounding within a given subgroup.
@@ A.4 Improving Finite-Sample Performance
and $\hat{e}(X_{i})$ is the estimated probability of being assigned to treatment, given the covariates $X$. $\hat{\tau}^{(g)}_{\textsc{w-adj}}$ provides a doubly robust approach to estimating the subgroup average treatment effect (e.g., Kang and Schafer, 2007), with potential efficiency gains in finite samples. Theorem 3.3 can be extended for $\hat{\tau}^{(g)}_{\textsc{w-adj}}$, allowing for valid inference and analogous hypothesis tests.
There are two components to $\hat{\tau}^{(g)}_{\textsc{w-adj}}$: (1) the covariate adjustment taking place within the regression, and (2) a re-weighting adjustment to balance the treatment and control groups. The covariate adjustment within the weighted regression allows researchers to leverage variation across covariates within a specific subgroup. In particular, because the subgroups are defined using covariates that are explanatory of the treatment effect variation, we generally expect a sparser set of covariates that explain the treatment effect variation in contrast to covariates that can explain variation in the outcome. As such, within each subgroup, while we expect relatively small amounts of treatment effect variation, there could still be residual variation in the outcomes that can be explained by the other covariates. The re-weighting component of $\hat{\tau}^{(g)}_{\textsc{w-adj}}$ ensures that within each subgroup, researchers can also account for finite-sample imbalances that are present (Xie et al., 2012).
@@
A.5 Detailed CDT Algorithm
@@ 5      Fit CATE model $\mathcal{M}$ using $\mathcal{D}^{train}$$\rightarrow\widehat{\mathcal{M}}$
6      Use $\widehat{\mathcal{M}}$ to estimate $\hat{\tau}(X_{i})$ for each unit $i$ in $\mathcal{D}^{train}$
@@ 12                Fit CATE model $\mathcal{M}$ using $\mathcal{D}^{train}_{in}$$\rightarrow\widehat{\mathcal{M}}$
13                Use $\widehat{\mathcal{M}}$ to estimate $\hat{\tau}^{(r)}(X_{i})$ for each unit $i$ in $\mathcal{D}^{train}_{out}$
@@ 14                Reverse roles of $\mathcal{D}^{train}_{in}$ and $\mathcal{D}^{train}_{out}$ and repeat lines 11-12
17               Average across cross-fits: $\hat{\tau}(X_{i})=\sum_{r=1}^{R}\hat{\tau}_{i}^{(r)}(X_{i})$ for each unit $i$ in $\mathcal{D}^{train}$
@@
21                Fit decision tree $m$ on $X^{train}$ to predict $\hat{\tau}(X_{i})$’s $\rightarrow$ estimated subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\ldots,\hat{\mathcal{G}}_{G}(X)\right\}$
22                // Honestly estimate subgroup average treatment effects
@@ 23                Estimate $\hat{\tau}^{(g)}$ via (6) for each $g=1,\ldots,G$ using $\mathcal{D}^{est}$
A.6 Additional Discussion of Teacher Model Selection Algorithm
The proposed teacher model selection procedure in Algorithm 1 requires choosing the desired tree depth $d$, or equivalently, the number of binary decision rules characterizing each subgroup. This ensures that there are the same number of subgroups (i.e., $2^{d}$) per tree (or partition). Having the same number of subgroups per tree is necessary to ensure that the subgroup Jaccard indices are comparable across both bootstraps and teacher models. Otherwise, a teacher model, such as a constant predictor, that results in fewer decision tree splits (or a smaller depth $d$) is inherently going to be more stable than a teacher model that results in more decision tree splits.
@@ A.7 Diagnostic Tools for CDT
Finally, in addition to evaluating the subgroup Jaccard stability across the different bootstrap samples in Algorithm 1, we can also track the selected features across these bootstraps for each chosen tree depth and plot their distribution (e.g., Figures A9-A14). Similar to the subgroup Jaccard stability, a more stable (or homogeneous) distribution of selected features should elucidate greater trust in that model or tree depth. On the flip side, a very heterogeneous distribution of selected features indicates that the estimated decision tree structure is highly volatile, depending heavily on the training data subset.
@@ Appendix B Proofs and Derivations
Then, applying the results from Escanciano (2020) and Banerjee and McKeague (2007):
@@
$=\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-2\mathbb{E}_{n}\left\{\hat{\tau}^{d}_{i}\cdot\hat{\tau}^{d}(X_{i},s)\right\}+\mathbb{E}_{n}\left\{\hat{\tau}^{d}(X_{i},s)^{2}\right\}-\left[\mathbb{E}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-2\mathbb{E}\left\{\hat{\tau}^{d}_{i}\cdot\tau^{d}(X_{i},s)\right\}+\mathbb{E}\left\{\tau^{d}(X_{i},s)^{2}\right\}\right],$
where $\hat{\tau}^{d}(X_{i},s):=\mathbbm{1}\{X_{i}\leq s\}\mathbb{E}_{n}(\hat{\tau}_{i}^{d}\mid X_{i}\leq s)+\mathbbm{1}\{X_{i}>s\}\mathbb{E}_{n}(\hat{\tau}_{i}^{d}\mid X_{i}>s\}$ is the sample analog to $\tau^{d}(X,s)$. Then:
$=\underbrace{\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-\mathbb{E}\left\{\left(\hat{\tau}_{i}^{d}\right)^{2}\right\}}_{(a)}-\underbrace{\left[\mathbb{E}_{n}\left\{\hat{\tau}^{d}(X_{i},s)^{2}\right\}-\mathbb{E}\left\{\tau^{d}(X_{i},s)^{2}\right\}\right]}_{(b)}.$
@@
$\mathbb{E}_{n}$$\left\{\hat{\tau}^{d}(X_{i};s)^{2}\right\}-\mathbb{E}\left\{\tau^{d}(X_{i};s)^{2}\right\}$
@@
$=$ $\left(\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\right)\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}+\left(\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]\right)\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}$
@@
To begin, we can take the expectation of $\hat{\tau}^{(g)}$ to show that it is a function of $\mathbb{E}(\tau_{i}\mid\hat{\mathcal{G}}_{g}(X_{i})=1)$.
$\mathbb{E}$ $\left(\hat{\tau}^{(g)}\right)=\mathbb{E}\left\{\frac{1}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}Z_{i}Y_{i}\hat{\mathcal{G}}_{g}(X_{i})-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}(1-Z_{i})Y_{i}\hat{\mathcal{G}}_{g}(X_{i})\right\}$
$=\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Z_{i}}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}\right\}Y_{i}(1)\hat{\mathcal{G}}_{g}(X_{i})-\mathbb{E}\left\{\frac{1-Z_{i}}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}\right\}Y_{i}(0)\hat{\mathcal{G}}_{g}(X_{i})$
@@
As such:
$\left|\mathbb{E}(\hat{\tau}^{(g)})-\tau^{(g)}\right|$ $=\left|\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\left\{Y_{i}(1)-Y_{i}(0)\right\}-\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\left\{Y_{i}(1)-Y_{i}(0)\right\}\right|$
$=\left\lvert\frac{1}{n_{g}}\sum_{i=1}^{n}\left\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\}\left\{Y_{i}(1)-Y_{i}(0)\right\}\right\rvert$
$\leq\frac{1}{n_{g}}\sum_{i=1}^{n}\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert\bigg{\lvert}Y_{i}(1)-Y_{i}(0)\bigg{\rvert},$
From Theorem 3.1, we know that $\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert\stackrel{{\scriptstyle p}}{{\rightarrow}}0$.
Part 2. Variance derivation. To derive the variance of the subgroup difference-in-means estimator, we apply results from Miratrix et al. (2013), who derives the variance of a post-stratification estimator. We can extend the results by treating each subgroup as a strata, where the subgroup difference-in-means estimator is analogous to the difference-in-means estimator across a specific strata. We provide the full derivation for completeness.
@@
$=\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left(\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}-\frac{1}{n_{g}}\right)Y_{i}(1)^{2}+\frac{1}{n_{g}}\sum_{i\neq j:\hat{\mathcal{G}}_{g}(X)=1}\left(\frac{1}{n_{g}-1}\left(1-\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}\right)-\frac{1}{n_{g}}\right)Y_{i}(1)Y_{j}(1)$
@@
$=\frac{\beta_{g}(1)}{n_{g}}\text{var}_{g}\left\{Y_{i}(1)\right\}+\frac{\beta_{g}(0)}{n_{g}}\text{var}_{g}\left\{Y_{i}(0)\right\}-\frac{1}{n_{g}}\left[\text{var}_{g}\left\{Y_{i}(1)\right\}+\text{var}_{g}\left\{Y_{i}(0)\right\}-2\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\right\}\right]$

### v2

@@ A.1 Pruning
As in the supervised learning setting, decision trees for causal effect estimation (including in the second stage of CDT) are often pruned to avoid overfitting. There are generally two types of pruning: pre-pruning and post-pruning. Pre-pruning refers to techniques or early stopping criteria that are used to stop growing the decision tree during the tree construction process. Common pre-pruning approaches include setting constraints on the maximum depth of the tree, the minimum number of samples per node, or the minimum information gain from making each split. Post-pruning refers to techniques used to remove uninformative nodes from the decision tree after it has already been constructed. For example, in the rpart R package (Therneau et al. 1997), cross-validation is performed to choose the best complexity parameter $\alpha$, where $\alpha$ measures the cost of adding another split to the tree, and all nodes with a cost greater than $\alpha$ are subsequently removed or pruned from the tree.
[…]
@@ A.2 Extended discussion on valid teacher models
Let $s$ and $s^{d}$ respectively be the optimal splits using the original individual-level treatment effects $\tau_{i}$ and predicted individual-level treatment effects $\hat{\tau}_{i}^{d}$, as defined in Assumption (3.2). Also, define $v_{i}:=\tau_{i}-\hat{\tau}_{i}^{d}$ for each $i=1,\ldots,n$. We can then write
@@
$\mathbb{E}\left[\left\lvert\hat{\mathcal{G}}_{1}(X_{i})-\mathcal{G}_{1}(X_{i})\right\rvert\right]\leq 2M\left\lvert\hat{s}_{n}-s^{d}\right\rvert+\varepsilon,$
where $\varepsilon=2M|s^{d}-s|$.
A.3 Inference for observational settings
While the main manuscript focused on experimental settings, we can similarly estimate the subgroup ATE in observational settings. However, we must account for potential confounding within a given subgroup.
@@ A.3 Inference for observational settings
[…]
@@
A.4 Extension to infinite population model
In the main manuscript, we focus on the finite-sample setting for considering inference for the subgroup difference-in-means estimator. In settings when researchers are interested in estimating the population-level subgroup ATE, if the sample is a random draw from this larger target population, then we can show that the subgroup difference-in-means estimator will be a consistent estimator. To begin, let $\mathcal{D}$ represent the finite-sample study data. Then:
$\mathbb{E}(\hat{\tau}^{(g)})$ $=\mathbb{E}\left\{\mathbb{E}\left(\hat{\tau}^{(g)}\mid\mathcal{D}\right)\right\}$
$=\mathbb{E}\left[\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\left\{Y_{i}(1)-Y_{i}(0)\right\}\right]$
$=\mathbb{E}\left[\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\cdot\left\{Y_{i}(1)-Y_{i}(0)\right\}\right]+\mathbb{E}\left[\frac{1}{n_{g}}\sum_{i=1}^{n}\left\{\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\}\cdot\left\{Y_{i}(1)-Y_{i}(0)\right\}\right]$
Let $K_{\tau}=\max_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left\{Y_{i}(1)-Y_{i}(0)\right\}$:
$\leq\tau^{(g)}+\frac{1}{n_{g}}\sum_{i=1}^{n}K_{\tau}\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert$
From Theorem 3.1, $\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert\stackrel{{\scriptstyle p}}{{\rightarrow}}0$. As such, $\hat{\tau}^{(g)}\stackrel{{\scriptstyle p}}{{\rightarrow}}\tau^{(g)}$.
The variance of the subgroup difference-in-means estimator would look similar to the variance expression in Theorem 3.3, but with the population-level parameters. Furthermore, when considering the population-level inference, the cross correlation of potential outcomes will disappear. As a result, while the variance estimator will be the same for both inference in a finite-sample setting and an infinite super-population setting, we generally expect more uncertainty associated with inference for the population-level subgroup ATE.
If the study is not a random sample from target population, then we must account for potential distribution shifts between the study sample and the target population. Being able to transport subgroups across different studies to different target populations would require stronger assumptions (see, for example, Zhang et al. 2024).
A.5 Non-parametric test of treatment effect heterogeneity
[…]
$H_{0}:\tau^{(1)}=\ldots=\tau^{(G)}=\tau,$ (7)
for all subgroups $g\in\{1,\ldots,G\}$, with the following estimator:
A.6 Detailed CDT Algorithm
@@ 5      Fit CATE model $\mathcal{M}$ using $\mathcal{D}^{train}$$\rightarrow\widehat{\mathcal{M}}$
6      Use $\widehat{\mathcal{M}}$ to estimate $\tau_{i}$ for each unit $i$ in $\mathcal{D}^{train}$ $\rightarrow\hat{\tau}_{i}^{d}$
@@ 12                Fit CATE model $\mathcal{M}$ using $\mathcal{D}^{train}_{in}$$\rightarrow\widehat{\mathcal{M}}$
13                Use $\widehat{\mathcal{M}}$ to estimate $\hat{\tau}^{d}_{i}$ for each unit $i$ in $\mathcal{D}^{train}_{out}$ $\rightarrow\hat{\tau}^{d}_{i,r}$
@@ 14                Reverse roles of $\mathcal{D}^{train}_{in}$ and $\mathcal{D}^{train}_{out}$ and repeat lines 11-12
17               Average across cross-fits: $\hat{\tau}^{d}_{i}=\sum_{r=1}^{R}\hat{\tau}^{d}_{i,r}$ for each unit $i$ in $\mathcal{D}^{train}$
@@
21                Fit decision tree $m$ on $X^{train}$ to predict $\hat{\tau}_{i}^{d}$’s $\rightarrow$ estimated subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\ldots,\hat{\mathcal{G}}_{G}(X)\right\}$
22                // Honestly estimate subgroup ATEs
@@ 23                Estimate $\hat{\tau}^{(g)}$ via (4) for each $g=1,\ldots,G$ using $\mathcal{D}^{est}$
Figure A1: Given a decision tree (A), its decision splits correspond to binary cuts that partition the covariate space $\mathcal{X}$, shown in (B). Furthermore, the decision splits leading to each terminal node map to the collection of rules that define a subgroup, shown in (C).
A.7 Discussion of Teacher Model Selection Algorithm
We formalize the teacher model selection algorithm in Algorithm 2.
1 for $l=1,\ldots,L$ do
2      Fit $\mathcal{M}^{l}$ on $\mathcal{D}^{train}$ to obtain estimates of the individual-level treatment effects $\hat{\tau}^{d,l}_{i}$ for each unit in $\mathcal{D}^{train}$ (via lines 2-16 in Algorithm 1)
3      for $b=1,\ldots,B$ do
4           for $t=1,2$ do
5                Get bootstrap sample of $\{(X_{i},\;\hat{\tau}^{d,l}_{i}):i\in\text{training unit}\}$
6                Fit decision tree $m$ on the bootstrapped $X_{i}$’s to predict the bootstrapped $\hat{\tau}^{d,l}_{i}$’s $\rightarrow$ estimated decision tree $\hat{m}^{(b_{t})}$
8
9                end for
11
12                end for
13
14                end for
Algorithm 2 Selecting the teacher model for CDT
Notably, the proposed teacher model selection procedure in Algorithm 2 requires choosing the desired tree depth $d$, or equivalently, the number of binary decision rules characterizing each subgroup. This ensures that there are the same number of subgroups (i.e., $2^{d}$) per tree (or partition). Having the same number of subgroups per tree is necessary to ensure that the Jaccard subgroup stability indices are comparable across both bootstraps and teacher models. Otherwise, a teacher model, such as a constant predictor, that results in fewer decision tree splits (or a smaller depth $d$) is inherently going to be more stable than a teacher model that results in more decision tree splits.
@@ A.8 Diagnostic Tools for CDT
Finally, in addition to evaluating the subgroup Jaccard stability across the different bootstrap samples in Algorithm 2, we can also track the selected features across these bootstraps for each chosen tree depth and plot their distribution (e.g., Figures A10-A15). Similar to the subgroup Jaccard stability, a more stable (or homogeneous) distribution of selected features should elucidate greater trust in that model or tree depth. On the flip side, a very heterogeneous distribution of selected features indicates that the estimated decision tree structure is highly volatile, depending heavily on the training data subset.
@@ Appendix B Proofs and Derivations
[…]
@@
$=\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-2\mathbb{E}_{n}\left\{\hat{\tau}^{d}_{i}\cdot\hat{f}^{d}_{\tau}(X;s^{\prime})\right\}+\mathbb{E}_{n}\left\{\hat{f}^{d}_{\tau}(X;s^{\prime})^{2}\right\}$
$\quad-\left[\mathbb{E}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-2\mathbb{E}\left\{\hat{\tau}^{d}_{i}\cdot f^{d}_{\tau}(X;s^{\prime})\right\}+\mathbb{E}\left\{f^{d}_{\tau}(X;s^{\prime})^{2}\right\}\right],$
where $\hat{f}^{d}_{\tau}(X;s^{\prime}):=\mathbbm{1}\{X_{i}\leq s\}\mathbb{E}_{n}(\hat{\tau}_{i}^{d}\mid X_{i}\leq s)+\mathbbm{1}\{X_{i}>s\}\mathbb{E}_{n}(\hat{\tau}_{i}^{d}\mid X_{i}>s\}$ is the sample analog to $\tau^{d}(X,s)$. Then:
$=\underbrace{\mathbb{E}_{n}\left\{\left(\hat{\tau}^{d}_{i}\right)^{2}\right\}-\mathbb{E}\left\{\left(\hat{\tau}_{i}^{d}\right)^{2}\right\}}_{(a)}-\underbrace{\left[\mathbb{E}_{n}\left\{\hat{f}^{d}_{\tau}(X;s^{\prime})^{2}\right\}-\mathbb{E}\left\{f^{d}_{\tau}(X;s^{\prime})^{2}\right\}\right]}_{(b)}.$
@@
$\mathbb{E}_{n}$$\left\{\hat{f}^{d}_{\tau}(X;s^{\prime})^{2}\right\}-\mathbb{E}\left\{f^{d}_{\tau}(X;s^{\prime})^{2}\right\}$
@@
$=$ $\left(\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}\leq s\right\}\right]\right)\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}\leq s)^{2}$
$+\left(\mathbb{E}_{n}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]-\mathbb{E}\left[\mathbbm{1}\left\{X_{i}>s\right\}\right]\right)\cdot\mathbb{E}_{n}(\hat{\tau}^{d}_{i}\mid X_{i}>s)^{2}$
@@
Part 2. Unbiasedness. To begin, we will show that $\hat{\tau}^{(g)}$ is unbiased for the within-sample subgroup ATE:
$\mathbb{E}$ $\left(\hat{\tau}^{(g)}\right)=\mathbb{E}\left\{\left.\frac{1}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}Z_{i}Y_{i}\hat{\mathcal{G}}_{g}(X_{i})-\frac{1}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}\sum_{i=1}^{n}(1-Z_{i})Y_{i}\hat{\mathcal{G}}_{g}(X_{i})~{}\right|\mathcal{D}\right\}$
$=\sum_{i=1}^{n}\mathbb{E}\left\{\left.\frac{Z_{i}}{\sum_{i=1}^{n}Z_{i}\hat{\mathcal{G}}_{g}(X_{i})}~{}\right|\mathcal{D}\right\}Y_{i}(1)\hat{\mathcal{G}}_{g}(X_{i})-\mathbb{E}\left\{\left.\frac{1-Z_{i}}{\sum_{i=1}^{n}(1-Z_{i})\hat{\mathcal{G}}_{g}(X_{i})}~{}\right|\mathcal{D}~{}\right\}Y_{i}(0)\hat{\mathcal{G}}_{g}(X_{i})$
@@
[…]
@@
$=\frac{1}{n_{g}}\sum_{i:\hat{\mathcal{G}}_{g}(X_{i})=1}\left(\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}-\frac{1}{n_{g}}\right)Y_{i}(1)^{2}+$
$~{}~{}~{}~{}\frac{1}{n_{g}}\sum_{i\neq j:\hat{\mathcal{G}}_{g}(X)=1}\left(\frac{1}{n_{g}-1}\left(1-\mathbb{E}\left\{\frac{1}{n_{g}(1)}\right\}\right)-\frac{1}{n_{g}}\right)Y_{i}(1)Y_{j}(1)$
@@
$=\frac{\beta_{g}(1)}{n_{g}}\text{var}_{g}\left\{Y_{i}(1)\right\}+\frac{\beta_{g}(0)}{n_{g}}\text{var}_{g}\left\{Y_{i}(0)\right\}$
$~{}~{}~{}~{}~{}-\frac{1}{n_{g}}\left[\text{var}_{g}\left\{Y_{i}(1)\right\}+\text{var}_{g}\left\{Y_{i}(0)\right\}-2\text{cov}_{g}\left\{Y_{i}(1),Y_{i}(0)\right\}\right]$